# Semi-automatic Calculations of Multi-loop Feynman Amplitudes with AmpRed

### v1

@@ 3.3 Differential equations
  • AlphaD[exp, x] gives the differentiation of exp with respect to x.
  • AlphaDES[I0, x] constructs differential equations for a list of Feynman integrals $I_{0}$. It returns $\{I,M\}$, with $I$ a list of master integrals and $M$ a matrix such that $\frac{\partial I}{\partial x}I=M\cdot I$.
    AlphaDES has the following options:
@@
  • AlphaAddScale[I, y, {i,j}] adds an auxiliary scale y to the parametric integral I through the replacement $x_{i}\to yx_{j}$. If the last argument is omitted, AlphaAddScale chooses the pair according to the algorithm describe in sec. 2.3.
@@
    "Braches" specifies the branches while crossing branch points. The option value should be a list of rules of the form x0 -> s, which indicates that a monomial $(x-x_{0})^{i}$ with $x<x0$ should be understood as $(x_{0}-x)^{i}e^{is\pi}$. By default, $s=-1$.
@@ 3.5 Interfaces
  • ARQgraf[i->o, n] generates the n-loop amplitudes (or diagrams) for the process i->o, with i and o the incoming and outgoing particles respectively. To use ARQgraf, a Fortran compiler (like gfortran) should be available. ARGgraf uses parallel tange2018gnu to parallelize computations (similarly for the other interfaces).
@@
    "Model" specifies the model file of QGRAF. The built-in models are "QED", "QCD", and "WL" (Wilson lines). In all these models, fermions are represented by "F_". In the model "QED", photons are represented by "V_". In the model "QCD", gluons are represented by "V_1_" and photons are represented by "V_2_". The model "QCD" is included in the model "WL". In the model "WL", a Wilson line in represented by "W_". The anti-particle of a particle, say, a fermion, is represented by "F_m". User-defined model files are also allowed. For user-defined model files, the option value of "Model" should either be the file name or the file context as a string. The file name should be wrapped with single quotes, for example, "’./model.mod’". Two model files can be combined by putting them in a single list. For example, "Model"->{"QCD","WL"}.
@@
    "ImportResults" specifies whether to import the final results or not. If
    "ImportResults"->False, the results will be written to some files, and the file names will be returned. For processes with a large number of diagrams, it is suggested to set "ImportResults"->False to save memory.
@@ 4 Summary and discussion
AmpRed uses the algorithms (method I and method II) described in ref. Chen:2019fzm to reduce tensor integrals. In method I, tensor integrals are directly parametrized are expressed in terms of parametric integrals with negative indices. The IBP identities for the resulting parametric integrals are equivalent to the traditional momentum-space IBP identities. In method II, tensor integrals are first parametrized and partially reduced by using the techniques of polynomial reduction based on eq. (24). Then the unreduced integrals are further reduced by combining symbolic rules with a Laporta-like algorithm.
@@
Reduction with method II is much more efficient than that with method I. Currently, this method is only implemented in a pure Mathematica code. While implementing the full method with another language other than Mathematica is far from trivial, it is expected to implement parts of it in the future, such as the symbolic rules. An interface to Kira is provided, but AmpRed only uses the user-defined system of Kira to solve IBP identities generated by eq. (12) by brute force, and the symbolic rules are not implemented. An observation is that even if we solve IBP identities by brute force, the efficiency is comparable with traditional methods while generating differential-equations systems.
AmpRed uses the method developed in ref. Chen:2023hmk to calculate master integrals iteratively. For real integrals, this method proves to be more efficient than other methods on the market in most circumstances. But for complex integrals, sometimes it becomes time-consuming to use this method. Thus, for complex integrals, it is suggested to calculate them by solving differential equations (by using the function DESSolveN) and calculate the boundary conditions (chosen in the Euclidean region) by using this method (with the function AlphaIntEvaluateN). It is expected to refine the method of analytic continuation in the future.
@@ Acknowledgements.
This work is supported by Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008).

### v2

@@ 3.3 Differential equations
  • AlphaD[exp, x] gives the differentiation of exp with respect to x. The option Method specifies the method to carry out the differentiation. The option value should be either 1 or 2, corresponding to eq. (31) and eq. (30) respectively. By default, Method->2.
  • AlphaDES[I0, x] constructs differential equations for a list of Feynman integrals $I_{0}$. It returns $\{I,M\}$, with $I$ a list of master integrals and $M$ a matrix such that $\frac{\partial I}{\partial x}I=M\cdot I$. AlphaDES uses AlphaD to carry out the differentiation.
    In addition to those that inherit from AlphaD, AlphaDES has the following options:
@@
  • AlphaAddScale[I, y, {i,j}] adds an auxiliary scale y to the parametric integral I through the replacement $x_{i}\to yx_{j}$. If the last argument is omitted, AlphaAddScale chooses the pair according to the algorithm described in sec. 2.3.
@@
    "Braches" specifies the branches while crossing branch points. The option value should be a list of rules of the form x0 -> s, which indicates that a monomial $(x-x_{0})^{i}$ with $x<x_{0}$ should be understood as $(x_{0}-x)^{i}e^{is\pi}$. By default, $s=-1$.
@@ 3.5 Interfaces
  • ARQgraf[i->o, n] generates the n-loop amplitudes (or diagrams) for the process i->o, with i and o the incoming and outgoing particles respectively. To use ARQgraf, a Fortran compiler (like gfortran) should be available. ARQgraf uses parallel tange2018gnu to parallelize computations (similarly for the other interfaces).
@@
    "Model" specifies the model file of QGRAF. The built-in models are "QED", "QCD", and "WL" (Wilson lines). In all these models, fermions are represented by "F_". In the model "QED", photons are represented by "V_". In the model "QCD", gluons are represented by "V_1_" and photons are represented by "V_2_". The model "QCD" is included in the model "WL". In the model "WL", a Wilson line is represented by "W_". The anti-particle of a particle, say, a fermion, is represented by "F_m". User-defined model files are also allowed. For user-defined model files, the option value of "Model" should either be the file name or the file context as a string. The file name should be wrapped with single quotes, for example, "’./model.mod’". Two model files can be combined by putting them in a single list. For example, "Model"->{"QCD","WL"}.
@@
    "ImportResults" specifies whether to import the final results or not. If "ImportResults"
    ->False
    , the results will be written to some files, and the file names will be returned. For processes with a large number of diagrams, it is suggested to set "ImportResults"
    ->False
    to save memory.
@@ 3.7 Examples
Some of these examples are described in detail in this section.
3.7.1 Reduction and numerical evaluation
In this subsection, we introduce functions to reduce integrals and numerically evaluate master integrals through a simple double-box example. We do the reduction by using both the built-in function AlphaReduce and the Kira interface.
The package can be loaded by running
In[16]:= AROptions={"UserKira"->True};
<<AmpRed‘
Before the calculations, we should set the kinematics first. That is, we express Lorentz invariants in terms of independent Mandelstam variables. And to parallelize the calculations, we launch some kernels.
In[17]:= SetKinematics[{p[1], p[2]} -> {k[1], k[2]}, {0, 0} -> {0, 0}];
LaunchKernels[];
Here the first argument of SetKinematics specifies the incoming and outgoing momenta, and the second argument specifies the corresponding masses. Here we set all the masses to 00. SetKinematics automatically sets Lorentz invariants in terms of Mandelstam variables $s$ and $t$. The symbols for Mandelstam variables can be changed by setting the option "Mandelstam". Alternatively, one can set the kinematics manually (See the introduction to Pair and SP in sec. 3.1.)
We define a standard tensor integral by using TI (See sec. 3.2).
In[18]:= int = TI[{l[1], l[2]}, {{k[1] + l[1]}, {k[1] + l[1] + l[2]},
{k[2] - l[1]}, {k[2] - l[1] - l[2]}, {k[2] - l[1] - l[2] - p[1]},
{l[1]}, {l[2]}}, FV[l[1], a]*FV[l[1], b]]
Out[18]= $-\frac{\pi^{-D}}{\Gamma(\epsilon+1)^{2}}\int d^{D}l_{1}d^{D}l_{2}$
$\frac{l_{1}^{a}l_{1}^{b}}{l_{1}^{2}l_{2}^{2}\left(k_{2}-l_{1}\right)^{2}\left(k_{1}+l_{1}\right)^{2}\left(k_{2}-l_{1}-l_{2}\right)^{2}\left(k_{1}+l_{1}+l_{2}\right)^{2}\left(k_{2}-l_{1}-l_{2}-p_{1}\right)^{2}}$
The integral reduction can be carried out by using the built-in function AlphaReduce.
In[19]:=res1=AlphaReduce[int];
AlphaReduce implements both method I and Method II described in sec. 2.2. The method can be specified through the option Method. By default, method II is used. Alternatively, the integrals can be reduced by using the Kira interface DoKira.
In[20]:=res2=DoKira[int];
Internally, DoKira uses AlphaParametrize to parameterize tensor integrals in terms of parametric integrals AlphaInt[...] and then reduces parametric integrals by using Kira. By default, DoKira chooses the reduction method automatically according to the structure of the expression. Specifically, if all integrals in the expression are standard loop integrals, DoKira uses the built-in reduction system of Kira to carry out the reduction. If the integrals are with momentum-space correspondences but non-standard (that is, integrals with delta and Heaviside theta functions), the reduction will be carried out by using the user-defined system of Kira, with the IBP identities generated by eq. (23). For integrals with no momentum-space correspondences, the reduction is carried out by the user-defined system with IBPs generated by eq. (12). One can also choose the method manually by setting the options "UserDefinedSystem" and Method.
The master integrals are numerically evaluated by using the function AlphaIntEvaluateN. In principle, one can directly apply AlphaIntEvaluateN to the original expression int. Nevertheless, in practical calculations, for a better performance, it is suggested to carry out the calculations step by step. Here, since int is already reduced to res1, we can evaluate res1 instead.
[…]
AlphaIntEvaluateN returns the numerical result of the expression, together with a list of numerical results of all master integrals. By default, AlphaIntEvaluateN uses AlphaReduce to carry out the integral reduction, which can be changed by setting the option AlphaReduce. For example, one can use DoKira to do the integral reduction by adding an option
AlphaReduce->DoKira.
3.7.2 Differential equations
In this subsection, we use a simple massive-banana example to show how to implement the differential equation method with AmpRed .
Again, we set the kinematics first.
In[22]:= SP[p, p] = x;
m = 1;
We use DoKira to do the integral reduction.
In[23]:= SetOptions[DoKira, {"UserDefinedSystem" -> True, Method -> 2}];
SetOptions[AlphaDES, {Method -> 2, AlphaReduce -> DoKira}];
The integral to be calculated is
In[24]:= int = TI[{l[1], l[2], l[3]}, {{l[1], m}, {l[2], m}, {l[3], m},
{l[1] + l[2] + l[3] + p, m}}]
Out[24]= $\frac{i\pi^{-3D/2}}{\Gamma(\epsilon+1)^{3}}\int d^{D}l_{1}d^{D}l_{2}d^{D}l_{3}\frac{1}{\left(l_{1}{}^{2}-1\right)\left(l_{2}{}^{2}-1\right)\left(l_{3}{}^{2}-1\right)\left(\left(l_{1}+l_{2}+l_{3}+p\right){}^{2}-1\right)}$
We use AlphaDES to construct the differential equation system.
In[25]:=des = AlphaDES[I*int, x];
We choose the boundary of the differential equations to be at $x=0$, which is in the Euclidean region. The boundary conditions of the differential-equation system can be determined by matching the asymptotic solutions of the differential-equation system with the asymptotic expansions of the master integrals. The former can be obtained by using DESAsymptoticSolve, and the latter can be obtained by using AlphaSeries. These operations are implemented in the function DESBoundary.
In[26]:=bnd = DESBoundary[des, x];
The boundary conditions are expressed in terms of parametric integrals, which can be numerically evaluated using AlphaIntEvaluateN.
In[27]:= bndn = AlphaIntEvaluateN[Last /@ bnd, 12, {}, PrecisionGoal -> 40,
AlphaReduce -> DoKira];
With the boundary conditions fixed, we can solve the differential equations numerically using the function DESSolvN. Notice that DESSolveN allows us to solve the differential equations at an arbitrary point, including the singularities. For example, we may solve the differential equations at the singularity $x=4$ by running
In[28]:= sol = DESSolveN[des[[2]], {x, 0, 4}, 8, MapThread[Rule, {First /@ bnd,
bndn[[1]]*eps^4}], PrecisionGoal -> 20]/eps^4;
Here the boundary constants are multiplied by a factor of eps^4 to make them free of poles in eps.
3.7.3 Calculation of a full amplitude
In this subsection, we demonstrate how to use AmpRed to calculate a full amplitude by calculating the two-loop correction to Higgs two-photon decay.
For simplicity, we set all Lorentz invariants to numbers.
In[29]:= LaunchKernels[];
SetKinematics[{P} -> {p[1], p[2]}, {FA["MH"]} -> {0, 0}];
FA["MH"] = FA["EL"] = 1;
FA["MT"] = 2;
FA["MW"] = 1/2;
FA["SW"] = 1/2;
We generate the amplitude using FeynArts, which is saved in the file
"AmpRed/examples/HiggsDecay_TwoLoop_Amplitude.m". We can directly import it and convert it to the format of AmpRed using the function FA2AR.
In[30]:= amp=FA2AR[Get["HiggsDecay_TwoLoop_Amplitude.m"],ToFeynmanInt->True]/.
{_Incoming->P,Outgoing->p}//SpinorChainSimplify//ColourSimplify;
Here the function SpinorChainSimplify calculates all the traces of Dirac gamma matrices, and ColourSimplify simplifies chains of color matrices. We use the built-in function AlphaReduce to reduce the integrals.
In[31]:=amp1 = AlphaReduce[amp];
We can check that the result satisfies the Ward identity.
In[32]:= Total[amp1]/.PolarizationVector[p[1],__]->p[1]//ContractLI//Simplify
Out[32]= 00 0
Finally we calculate the master integrals numerically by
In[33]:=amp2 = AlphaIntEvaluateN[amp1/.ColourN["F"]->3, 8, {}, PrecisionGoal->20];
@@ 4 Summary and discussion
AmpRed uses the algorithms (method I and method II) described in ref. Chen:2019fzm to reduce tensor integrals. In method I, tensor integrals are directly parametrized and are expressed in terms of parametric integrals with negative indices. The IBP identities for the resulting parametric integrals are equivalent to the traditional momentum-space IBP identities. In method II, tensor integrals are first parametrized and partially reduced by using the techniques of polynomial reduction based on eq. (24). Then the unreduced integrals are further reduced by combining symbolic rules with a Laporta-like algorithm.
@@
Reduction with method II is much more efficient than that with method I. Currently, this method is only implemented in a pure Mathematica code. While implementing the full method with another language other than Mathematica is far from trivial, parts of it are expected to be implemented in the future, such as the symbolic rules. An interface to Kira is provided, but AmpRed only uses the user-defined system of Kira to solve IBP identities generated by eq. (12) by brute force, and the symbolic rules are not implemented. An observation is that even if we solve IBP identities by brute force, the efficiency is comparable with traditional methods while generating differential-equation systems.
AmpRed uses the method developed in ref. Chen:2023hmk to calculate master integrals iteratively. For real integrals, this method proves to be more efficient than other methods on the market in most circumstances. However, for complex integrals, it sometimes becomes time-consuming to use this method. Thus, for complex integrals, it is suggested to calculate them by solving differential equations (by using the function DESSolveN) and calculate the boundary conditions (chosen in the Euclidean region) by using this method (with the function AlphaIntEvaluateN). It is expected to refine the method of analytic continuation in the future.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NNSFC) under Grant No. 12405095 and Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008).